OpenStudy (anonymous):

multiply the following: 234x.0987454/234=

Parth (parthkohli):

234 gets cancelled actually.

Parth (parthkohli):

That leaves 0.0987454

OpenStudy (ash2326):

\[234\times \frac {0.0987454} {234} \]

OpenStudy (ash2326):

\[\cancel {234}\times \frac {0.987454} {\cancel {234}}\]

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow x \times {y \over x} = \cancel{x} \times {y \over \cancel{x}} }\)

OpenStudy (anonymous):

why do you divide 0.0987454 by 10 after the 234's cancel out leaving 0.987454

Parth (parthkohli):

Maybe it's a mistake from ash hehe

OpenStudy (anonymous):

Ahh ok :)

OpenStudy (anonymous):

so the answer is .0987454?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok that wa easy. lol

Parth (parthkohli):

Exactly

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[80\sqrt{32}\]

OpenStudy (anonymous):

what is\[80/\sqrt{32}\]

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \sqrt{32} = \sqrt{2 * 2 * 2 * 2 * 2} = 4\sqrt{2} }\)

Parth (parthkohli):

What do you multiply to 4 root 2 to get 80?

OpenStudy (anonymous):

20

Parth (parthkohli):

No...then you will get 80 root 2

OpenStudy (anonymous):

im lost :(

OpenStudy (anonymous):

5

Parth (parthkohli):

no no

Parth (parthkohli):

sqrt 2 to get 2, then further multiply 20. So, 20 sqrt 2.

OpenStudy (anonymous):

yeah had said 20. i know sqrt stays the same

OpenStudy (anonymous):